OpenStudy (anonymous):

solve for W: p= 2L+2w

jimthompson5910 (jim_thompson5910):

p= 2L+2w p- 2L = 2w 2w = p - 2L what's next?

OpenStudy (anonymous):

do i multiply or divide by 2?

jimthompson5910 (jim_thompson5910):

divide both sides by 2

jimthompson5910 (jim_thompson5910):

to isolate w

jimthompson5910 (jim_thompson5910):

because you want to undo multiplication by 2

OpenStudy (anonymous):

so the answer would be w=p-2L/2?

jimthompson5910 (jim_thompson5910):

yes \[\Large w = \frac{p-2L}{2}\]

OpenStudy (anonymous):

okay thank you! wish you were my teacher!!

jimthompson5910 (jim_thompson5910):

yw